# Few-shot Class-incremental Learning: A Survey

### v2

@@ 6.1 Human-machine Gap in FSCIL
Currently, some IL research, such as the method proposed by ZKudithipudi et al. [113] that emulates the Drosophila’s mushroom body’s mechanisms, attempts to enhance model memory capabilities by bio-inspired intelligence. However, a systematic bio-inspired approach in FSCIL is yet to be established. Current FSCIL models lack in associative learning and abstract thinking in limited sample learning, and there’s room for improvement in prior knowledge acquisition. These models typically use one model for storing all knowledge from continual learning, suggesting the need for exploring multi-modular knowledge storage and long-short term memory mechanisms. Additionally, FSCIL requires proactive strategies for knowledge consolidation, updating, and personalized management, such as actively forgetting infrequent knowledge, reinforcing challenging knowledge, and integrating consistent knowledge.
@@ 6.3 Knowledge Acquisition and Update in FSCIL
Base Stage: Effective initialization of the backbone is crucial for ensuring base class performance and generalization for future incremental classes. Current methods often rely on a large number of base class samples for backbone network initialization, which may not align with reality and whose generalization capabilities are difficult to accurately assess. To enhance generalization, researchers have tried introducing strategies like self-supervised learning and forward compatibility, but these usually depend on sufficient base class data. There is a lack of research on initial knowledge acquisition without specific requirements for the base data. Therefore, exploring methods to enrich initial stage knowledge acquisition is important. From the data perspective, increasing data diversity and improving knowledge learning strategies, such as exploring data augmentation, data generation, introducing unsupervised data, and optimizing backbone learning methods, are essential. Additionally, introducing pre-trained models and other prior knowledge can be considered. For instance, foundation models like CLIP, SAM, and GPT, which combine self-supervised or semi-supervised pre-training with prompt engineering, have shown excellent generalization and transfer capabilities, offering new possibilities for enhancing FSCIL model performance.
Incremental Stage: In incremental sessions, models typically initialize with weights from prior phases, focusing on learning new classes and preserving existing knowledge. Challenges arise from limited new samples and restricted access to complete old data, making effective learning of new categories and old knowledge retention pivotal. Current solutions include freezing the backbone network and using class prototype averaging, demanding robust generalization and discrimination from the network, yet possibly leading to reduced performance as new classes increase. An alternative is maintaining key parameters for new class learning, though this risks diminishing old class performance and complicates parameter evaluation due to deep learning models’ opaque nature. KD is also commonly used, but how to effectively learn new category and select efficient old samples for distillation is still a direction to be further explored.
@@ 6.4 Applications and Safety in FSCIL
This section primarily discusses the potential application scenarios of FSCIL and the privacy concerns associated with its applications.
Application Scenarios: Current FSCIL research mainly targets image classification, with emerging yet non-systematic studies in visual object detection, natural language processing, lip reading, remote sensing, and robotics. Most work evaluates performance on benchmark datasets, with real-world applications still evolving. In many application scenarios, the demand for few-shot continuous learning capability is significant. For instance, in applications like video analysis, service robotics in hotels, and autonomous driving, the need for FSCIL technology is evident. These fields often require learning new classes from limited data, maintaining high accuracy with scarce samples, and adapting to new categories in dynamic environments, underscoring FSCIL’s importance and potential.

### v3

@@ 6.1 Human-machine Gap in FSCIL
Currently, some IL research, such as the method proposed by ZKudithipudi et al. [115] that emulates the Drosophila’s mushroom body’s mechanisms, attempts to enhance model memory capabilities by bio-inspired intelligence. However, a systematic bio-inspired approach in FSCIL is still lacking. Current FSCIL models lack associative learning and abstract thinking in limited sample learning, and there is room for improvement in prior knowledge acquisition. These models typically use one model for storing all knowledge from continual learning, suggesting the need for exploring multi-modular knowledge storage and long-short term memory mechanisms. Additionally, FSCIL requires proactive strategies for knowledge consolidation, updating, and personalized management, such as actively forgetting infrequent knowledge, reinforcing challenging knowledge, and integrating consistent knowledge.
@@ 6.3 Knowledge Acquisition and Update in FSCIL
Base Stage: Effective initialization of the backbone is crucial for ensuring base class performance and generalization for future incremental classes. Current methods often rely on a large number of base class samples for backbone network initialization, which may not align with reality and whose generalization capabilities are difficult to accurately assess. To enhance generalization, researchers have tried introducing strategies like self-supervised learning and forward compatibility, but these usually depend on sufficient base class data. There is a lack of research on initial knowledge acquisition without specific requirements for the base data. Therefore, exploring methods to enrich initial stage knowledge acquisition is important. From the data perspective, increasing data diversity and improving knowledge learning strategies, such as exploring data augmentation, data generation, introducing unsupervised data, and optimizing backbone learning methods, are essential. Additionally, introducing pre-trained models and other prior knowledge can be considered. For instance, foundation models like CLIP, SAM, and GPT, which combine self-supervised or semi-supervised pre-training with prompt engineering, have shown excellent generalization and transfer capabilities, offering new possibilities for enhancing FSCIL model performance. Some recent works have attempted to incorporate foundation models to address FSCIL challenges. For example, D’Alessandro et al. [116] designed a prompt learning strategy for CLIP in FSCIL, and Zhang et al. [117] leveraged the RETFound foundation model to enhance feature learning in few-shot class-incremental retinal disease recognition. However, these approaches have not yet become mainstream, and attention to fairness in experimental comparisons is still necessary.
Incremental Stage: In incremental sessions, models typically initialize with weights from prior phases, focusing on learning new classes and preserving existing knowledge. Challenges arise from limited new samples and restricted access to complete old data, making effective learning of new categories and old knowledge retention pivotal. Current solutions include freezing the backbone network and using class prototype averaging, demanding robust generalization and discrimination from the network, yet possibly leading to reduced performance as new classes increase. An alternative is maintaining key parameters for new class learning, though this risks diminishing old class performance and complicates parameter evaluation due to deep learning models’ opaque nature. KD is also commonly used, but how to effectively learn new categories and select efficient old samples for distillation is still a direction to be further explored.
@@ 6.4 Applications and Safety in FSCIL
Application Scenarios: Current FSCIL research mainly targets image classification, with emerging yet non-systematic studies in visual object detection, natural language processing, lip reading, remote sensing, and robotics. Most work evaluates performance on benchmark datasets, with real-world applications still evolving. In many application scenarios, the demand for the few-shot continuous learning capability is significant. For instance, in applications like video analysis, service robotics in hotels, and autonomous driving, the need for FSCIL technology is evident. These fields often require learning new classes from limited data, maintaining high accuracy with scarce samples, and adapting to new categories in dynamic environments, underscoring FSCIL’s importance and potential.